Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability

Expert replies
by AbhiS » Tue Nov 12, 2013 4:42 am
Each day after an item is lost the probability of finding that item is halved. If 3 days after a certain item is lost the probability of finding it has dropped to 1/64, what was the initial probability of finding the item?

a)1/32
b)1/8
c)1/4
d)1/2
e)1

[spoiler]
OA - D OR B, I have 2 sources who have sighted 2 different OA for the same question.
However I feel the OA should be D

This is how i reached to the D which i feel is correct because it mentions that - Each day after an item is lost the probability of finding that item is halved

i used the formula - (x)*(1/2*x)*(1/4*x) and solved; simplified solution (1/2)*(1/2*1/2)*(1/2*1/2*1/2) = 1/64

The other explanation was -

x * 1/2*1/2*1/2 = 1/64
Hence x = 1/8[/spoiler]
Join the discussion
Source: — Problem Solving |

by Mathsbuddy » Tue Nov 12, 2013 4:54 am
Working forwards in time:
P x 1/2 x 1/2 x 1/2 = 1/64

Simplified this is:
P/8 = 1/64

Therefore P = 6/64 = 1/8 (ANSWER B)

Alternatively, we could work backwards in time from 1/64:
1/64 x 2 x 2 x 2 = P

Again P = 1/8
Join the discussion

by theCodeToGMAT » Tue Nov 12, 2013 4:55 am
I believe it's {B}

let initial probability = x
1st day = x/2
2nd day = x/4
3rd day = x/8 -(1)

Given: probability after 3 days = 1/64 -(2)

equation (1) & (2)

x/8 = 1/64

x = 1/8

Answer [spoiler]{B}[/spoiler]
R A H U L
Join the discussion

by Mathsbuddy » Tue Nov 12, 2013 5:23 am
AbhiS wrote:Each day after an item is lost the probability of finding that item is halved. If 3 days after a certain item is lost the probability of finding it has dropped to 1/64, what was the initial probability of finding the item?

a)1/32
b)1/8
c)1/4
d)1/2
e)1

[spoiler]
OA - D OR B, I have 2 sources who have sighted 2 different OA for the same question.
However I feel the OA should be D

This is how i reached to the D which i feel is correct because it mentions that - Each day after an item is lost the probability of finding that item is halved

i used the formula - (x)*(1/2*x)*(1/4*x) and solved; simplified solution (1/2)*(1/2*1/2)*(1/2*1/2*1/2) = 1/64

The other explanation was -

x * 1/2*1/2*1/2 = 1/64
Hence x = 1/8[/spoiler]

The second method was correct (x = 1/8)
The reason the other method is wrong is that the formula has "x" repeated:
(x)*(1/2*x)*(1/4*x) = 1/64
The probability is not multiplied by half the probability; it is only multiplied by half.
So the formula should just be:
(x)*(1/2)*(1/2)*(1/2) = 1/64
(1/8)*(1/2)*(1/2)*(1/2) = 1/64
Therefore x = 1/8


The error of also multiplying by "x" each day is akin to saying
P(Day 0) and P(Day 1) and P(Day 2) and P(Day 3) = 1/64
which has no justification in this context.

In other words:
Day 0: P = 1/8
Day 1: P1 = 1/8 x 1/2 = 1/16
Day 2: P2 = 1/8 x 1/2 x 1/2 = 1/32
Day 3: P3 = 1/8 x 1/2 x 1/2 x 1/2 = 1/64

A shorthand formula would use powers (indices), where "^" denotes to the power of.
(Starting probability) x (fractional daily change)^(no. of days) = (final probability)

P x (1/2)^3 = 1/64

So, P/8 = 1/64 -> P = 1/8.

This formula is the principle method of calculating compound interest on money invested, for example:

(Starting Money) x (100% + annual interest rate%)^(Number of years) = (Final sum)

Note that care must be taken when using such formulae; understand whether the scenario involves starting the interest calculations on year 0 or year 1, and adjust accordingly.
Join the discussion

by GMATGuruNY » Tue Nov 12, 2013 6:08 am
AbhiS wrote:Each day after an item is lost the probability of finding that item is halved. If 3 days after a certain item is lost the probability of finding it has dropped to 1/64, what was the initial probability of finding the item?

a)1/32
b)1/8
c)1/4
d)1/2
e)1
We can PLUG IN THE ANSWERS, which represent the initial probability.
Each day, the probability is multiplied by 1/2.
The correct answer choice will yield P = 1/64 on the 3rd day.
Answer choice A is way too small, since it implies that P = 1/2 * 1/32 = 1/64 after only one day.
Eliminate A.

Answer choice C: 1/4
1st day: P = 1/2 * 1/4 = 1/8.
2nd day: P = 1/2 * 1/8 = 1/16.
3rd day: P = 1/2 * 1/16 = 1/32.
The resulting probability is too great, implying that the correct answer choice must be SMALLER.

The correct answer is B.
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion